The square of Mark's age 3 years ago is 6 times the age he will be in 9 years. What is his age now?
Answer by lwsshak3(11628)

The square of Mark's age 3 years ago is 6 times the age he will be in 9 years. What is his age now?
..
let x=Mark's age now
3 years ago,
Mark's age=(x-3)
9 years from now,
Mark's age=(x+9)
(x-3)^2=6(x+9)
x^2-6x+9=6x+54
x^2-12x-45=0
(x-15)(x+3)=0
x+3=0
x=-3 (reject)
or
x-15=0
x=15 (ans)